What is the inventory conversion period for O'Brian's if it has sales of $320,000,

an average inventory of $5,333 and a cash conversion cycle of 20 days?

Assume the cost of sales is 55% of sales.


1
Expert's answer
2021-08-04T13:58:02-0400

The inventory conversion period is based on the concept of days of inventory take into sales. It can be calculated by dividing the total days of the year by the inventory turnover ratio.


Inventory conversion period = 365 / Inventory turnover ratio


Inventory turnover ratio = Cost of goods sold / Average inventory 


Cost of goods sold = Sales * Percentage of cost of soles


= 320,000 * 55 % = 176,000


Inventory turnover ratio = 176000 / 5,333 = 33 times


Inventory conversion period = 365 / 33 = 11 Days